OpenStudy (anonymous):

solve the trigonometric equation... tan(x+π) + 2sin(x+π)=0

OpenStudy (anonymous):

in which interval?

OpenStudy (anonymous):

just general

OpenStudy (anonymous):

u=x+π tanu+2sinu=0 (sinu/cosu)+2sinu=0 ((2cosu+1)sinu)/cosu)=0 (2cosu+1)sinu=0 2cosu+1=0 ____ cosu=-0.5 u=(2/3)π+2kπ or u=(-2/3)π+2kπ or sinu=0 u=kπ

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

welcome